SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

   
Prev Next


Date: March 06, 2008
Shift Taker: David Doll
-----------
General Features:
(theta,phi)

Warm crystals in hitmap	(52,63), (56,63), (55,64)
Warm crystals in 0.01 GeV hitmap (17,43), (30,45), (32,46), (36,64)
Warm crystals in 0.1 GeV hitmap (30,45), (32,46), (36,64)
Warm crystals in 1 GeV hitmap      

Cool crystal  in hitmap          
Cool crystal  in 0.1 GeV hitmap (29,84) 
Cool crystal  in 1 GeV hitmap    		

0080218: [1]
		TC Damage: OOO, OOS ~10^-3 (0-4) all channels
0080219: [1]
0080220: [1]
		TC Damage: OOO, OOS ~5x10^-3 (0-4) all channels
0080221: [1]
		[2]
0080222: [1]
0080223: [1]
0080224: [1]
                TC Damage: OOO, OOS ~10^-3 (0-4, 4-9) all channels
0080225: [1], [2]
                TC Damage: OOO, OOS ~2x10^-3 (0-4) all channels
0080226: [1]
0080227: [1]
0080228: warm channel at (17,43) in no cuts hitmap
		[1]
0080229: warm channel at (17,43) in no cuts hitmap
0080230: warm channel at (17,43) in no cuts hitmap
		trough very low in E=0.01 proj. plot.
                TC Damage: OOO, OOS ~2x10^-3 (0-4) all channels
0080231: warm channel at (17,43) in no cuts hitmap
0080232: warm channel at (17,43) in no cuts hitmap
0080233: warm channel at (17,43) in no cuts hitmap
0080234: warm channel at (17,43) in no cuts hitmap
		TC Damage: OOS ~10^-3 (0-4) all channels.
0080235: TC Damage: OOO, OOS ~2x10^-3 (0-4) all channels
0080236: warm channel at (17,43) in no cuts hitmap 
0080237: warm channel at (17,43) in no cuts hitmap
0080238: warm channel at (17,43) in no cuts hitmap
0080239: warm channel at (17,43) in no cuts hitmap
0080240: 2nd peak around 200 in projection plot (no cuts)
0080241: 2nd peak around 200 in projection plot (no cuts)
0080242: warm channel at (17,43) in no cuts hitmap
0080243: warm channel at (17,43) in no cuts hitmap
0080244: warm channel at (17,43) in no cuts hitmap
0080245: warm channel at (17,43) in no cuts hitmap
                TC Damage: OOO, OOS ~10^-3 (0-4) all channels.
0080246: warm channel at (17,43) in no cuts hitmap
0080247: everythings fine
0080248: warm channel at (17,43) in no cuts hitmap
                TC Damage: OOO, OOS ~10^-3 (0-4) all channels.
0080249: warm channel at (17,43) in no cuts hitmap
                TC Damage: OOO, OOS ~10^-3 (0-4) all channels.
0080250: warm channel at (17,43) in no cuts hitmap
		2nd peak above 2000 in no cuts projection plot
0080251: warm channel at (17,43) in no cuts hitmap
0080252: warm channel at (17,43) in no cuts hitmap 

	[1]= 	No trough structure on left side of E=0.01 projection plot.
	[2]=	Large tail on the right side of the no cuts and E=0.01 Digi Multiplicity plot
        [3]=    No fall off on left side of phi=30 peak in the energy profile.
                No fall off on right side of phi=90 peak in the energy profile.

Crystal Spectra:
Badness 1 (37, 43),(39-40,43)
Badness 2 (38, 43)

Lightpulser:
4 channels dead, 0 channels out of time. Last updated on: 04-March 2008 19:04 PST


 

  Links

  • Shift Report Index
  • Back to the Shifter instructions